Find the curvature, the radius and the center of curvature at a point.


x=t, y=1/t ; t=1


1
Expert's answer
2021-12-06T12:11:57-0500

curvature:

"K=\frac{|x'y''-y'x''|}{((x')^2+(y')^2)^{3/2}}"


"x'=1,x''=0"

"y'=-1\/t^2,y''=2\/t^3"


"K=\frac{2}{(1+1)^{3/2}}=\frac{1}{\sqrt 2}"


radius of curvature:

"R=1/K=\sqrt 2"


 center of curvature:

"x_c=x_0+Rsin|\theta|"

"y_c=y_0+Rcos|\theta|"

where "tan\theta" is slope of tangent

"tan\theta=f'(x_0)"

"x_0=1,y_0=1"

"f'(x)=y'\/x'=-1\/t^2"

"f'(x_0)=-1"

"|\theta|=45\degree"


"x_c=1+1=2"

"y_c=1+1=2"